OpenStudy (anonymous):

wingspan can run the 440 yard dash in 55 seconds, and Jack can run it in 88 seconds. How great a handicap must wingspan give Jack for the boys to finish the race at the same time? A.165 yards B.242 yards C.275 yards

OpenStudy (amistre64):

when: 440/55 = d speed; 440/88 = j speed j(55) + hs = 440 hs = 440/(440/50) maybe?

OpenStudy (amistre64):

hs = 440/(j*55)

OpenStudy (anonymous):

so a, b, or c?

OpenStudy (amistre64):

440 - j(55) looks better

OpenStudy (amistre64):

165 if i did it right :)